OpenStudy (nthenic_oftime):

Find the vertical asymptotes, if any, of the graph of the rational function. Show your work. f(x) = (x-4)/(x(x-4))

OpenStudy (welshfella):

there are asymptotes at the values of x which make the denominator zero.

OpenStudy (welshfella):

so basically you need to solve x(x - 4) = 0 the vertical asymptotes will be x = 'the value of a root'

OpenStudy (welshfella):

x(x - 4) = 0 means x = 0 or x-4 = 0 can you continue?

OpenStudy (nthenic_oftime):

how do i find the root that the value of x will be i understand that the denominator need be zero

OpenStudy (welshfella):

well one root is 0 so there an asymptote at x = 0 and if x - 4 = 0 what is x?

OpenStudy (welshfella):

I'm sure you can solve x - 4 = 0 ?!!

OpenStudy (welshfella):

add 4 to both sides x - 4 + 4 = 0 = 4

OpenStudy (welshfella):

thats 0 + 4 on the right

OpenStudy (welshfella):

sorry gotta go

OpenStudy (nthenic_oftime):

sorry i got distracted thank you :)

OpenStudy (nthenic_oftime):

yeah i can solve that lol

OpenStudy (welshfella):

ok